calculate how much stock solution and how much water are needed to prepare 375 ml of a 38 ug/ml solution given a stock solution of 2mg/m
1
Expert's answer
2020-09-30T07:57:46-0400

2mg/ml=2*10^(-3) g/ml

To prepare 375 ml 38microgram/ml or 38*10^(-6)g/ml solution from the given stock solution we use the formula

V1S1=V2S2

V1= final volume of solution=375ml

S1=final concentration of solution=38*10^-6 g/ml

V2= required volume of stock solution

S2= concentration of stock solution=2*10^-3g/ml

Putting all these value in above equation we get V2=V1S1/S2=7.125ml

And (375-7.125)=367.875 ml water required